### v2

# Evidence of anisotropic three-dimensional weak-localization in TiSe2 nanoflakes

Abstract
TiSe2 is a typical transition-metal dichalcogenide known for its charge-density wave order. In this study, we report the observation of an unusual anisotropic negative magnetoresistance in exfoliated TiSe2 nanoflakes at low temperatures. Unlike the negative magnetoresistance reported in most other transition-metal dichalcogenides, our results cannot be explained by either the conventional two-dimensional weak localization effect or the Kondo effect. A comprehensive analysis of the data suggests that the observed anisotropic negative magnetoresistance in TiSe2 flakes is most likely caused by the three-dimensional weak localization effect.
I Introduction
The transition metal dichalcogenide (TMD) TiSe2 has attracted considerable research attention over the past decades due to its fascinating feature of charge-density-wave (CDW) order [1, 2]. Similar to other layered TMD compounds, the TiSe2 crystal composes of triangular Se-Ti-Se van der Waals layers stacked along the axis. The CDW transition of TiSe2 occurs at K, below which a commensurate superlattice emerges [1, 2]. Although the exact origin of the CDW transition remains controversial, there has been a consensus that it is not due to Fermi surface nesting [3, 4, 5]. Furthermore, a long-standing debate persists regarding whether TiSe2 is a semimetal [6, 7, 8] or a narrow-gap semiconductor [9, 10, 11, 12, 13]. Recent results from angle-resolved photoemission spectroscopy (ARPES) and transport measurements indicate that the CDW state of TiSe2 is a semiconductor with a narrow band gap [10, 11, 12, 13]. Meanwhile, whether TiSe2 has a finite normal-state energy gap is still under discussion [7, 8, 9, 10].
Compared to studies on the CDW order in TiSe2, experiments focusing on the coherent transport processes have been relatively fewer. Recently, an out-of-plane negative magnetoresistance (NMR) was reported at low temperatures in both bulk samples [12] and exfoliated flakes [14] of TiSe2. In the latter case, the observed NMR was attributed to the two-dimensional weak-localization (2D WL) effect [14]. However, due to the lack of magnetoresistance (MR) data measured at sufficiently low temperatures and under in-plane magnetic fields, further investigations are still needed to fully elucidate the origin of the NMR observed in TiSe2.
In this article, we present our study of the magnetotransport properties of exfoliated TiSe2 nanoflakes. The MR data of TiSe2 flakes were measured down to mK in a 3He cryostat. Surprisingly, an unusual anisotropic NMR was observed in both perpendicular and in-plane magnetic fields. It was found that the NMR curves measured under in-plane fields were significantly broader than those measured under perpendicular fields, indicating a strong anisotropy of the NMR. Such an unusual anisotropic NMR deviates substantially from the theoretical expectations of both the 2D WL and Kondo effects, which are commonly associated with NMR at low temperatures. A detailed analysis of the data suggests that the observed anisotropic NMR in TiSe2 nanoflakes presumably originates from the 3D WL effect.
II Experimental Details
II.1 Crystal Growth
The TiSe2 single crystals used in this work were grown by chemical vapor transport. High-purity titanium powder (99.9%, Alfa Aesar) and selenium shots (99.99%, Alfa Aesar) were used as source materials. Before the growth, the source materials were pre-treated to remove any potential oxide layers on the surfaces. The pretreatment includes two steps: hydrogen reduction and sublimation recrystallization. Afterward, the refined titanium and selenium were sealed in a 10-cm-long evacuated quartz tube, with an atomic ratio Ti:Se=1:2.1. During the growth, the temperatures at the cold and hot ends of the quartz tube were optimized to 570 oC and 640 oC, respectively. The growth took approximately one week. After the growth, TiSe2 crystals with shiny facets were obtained, as shown in the inset of Fig. 1(a). The structure of the obtained TiSe2 crystals was confirmed by X-ray diffraction (XRD). Fig. 1(a) plots the XRD pattern of a typical TiSe2 crystal, in which sharp Bragg peaks of the (00) surfaces are clearly resolved, indicating a high crystallinity of the sample.
II.2 Device Fabrication and Transport Measurements
Thin TiSe2 flakes were exfoliated onto Si(0.5 mm)/SiO2(300 nm) substrates using Scotch tape. To prevent oxidation, the exfoliation was performed in a glove box filled with N2. After the exfoliation, flakes with regular shapes and flat surfaces were selected out for device fabrication. The Ti (5 nm)/Au (200 nm) electrodes were fabricated onto TiSe2 flakes using electron-beam lithography and electron-beam evaporation. To reduce contact resistance, the contact areas were cleaned in situ with Ar plasma prior to metal deposition. The data presented in this paper were obtained in two typical devices, labeled as devices A and B. The thickness of the TiSe2 flakes in devices A and B was measured to be 119 nm and 90 nm, respectively, using an atom force microscope (AFM).
The longitudinal resistivity and Hall resistivity of TiSe2 flakes were measured in a Hall-bar geometry using lock-in amplifiers, with an AC excitation current of 3 A at 17.78 Hz. No obvious overheating caused by excitation current was observed [15]. The measurements were performed in a 3He cryostat with a base temperature of 260 mK.
III Results and discussion
III.1 Temperature Dependence of Resistivity
The curves of TiSe2 flakes behave similarly to those of bulk TiSe2 crystals [1]. As plotted in Fig. 1(b), upon cooling down from room temperature, first increases until reaches a maximum at K. It then decreases with further lowering of temperature, exhibiting a metallic behavior. The most prominent feature of the curves is the anomalous resistivity peak centered around K, which has long been a focal point in previous studies. Early work speculated that the resistivity peak might be related to the excitons in the CDW ordering [16, 17, 18]. Recently, Watson et al. proposed that it could be explained by a simple two-band model [19], without invoking the CDW order in TiSe2.
Similar to bulk crystals, TiSe2 flakes also undergo a CDW transition during cooling. As shown in Fig. 1(c), the transition temperatures for devices A and B were determined to be 190 K and 194 K, respectively, based on the peaks in the second derivative of the curves. These values are in good agreement with the previously reported in bulk crystals [1] and exfoliated flakes [14] of TiSe2 .
Fig. 1(d) shows the low-temperature part of the curves normalized to the value at K. At zero magnetic field, the resistivity of TiSe2 flakes does not saturate at low temperatures but instead increases slightly with decreasing , leading to a resistivity upturn at K. The low-temperature resistivity upturns are completely suppressed by an in-plane field T.
Possible mechanisms responsible for such low-temperature resistivity upturns include WL, the Kondo effect, and electron-electron interactions (EEI). Among them, WL and the Kondo effect are sensitive to magnetic fields, whereas EEI is not [20]. Therefore, the suppression of the resistivity upturns by external fields rules out EEI as their cause.
III.2 Magnetotransport
The MR curves of devices A and B were measured in both perpendicular and in-plane magnetic fields at temperatures ranging from 0.26 K to 20.6 K, as plotted in Fig. 2. All curves exhibit distinct NMR in the low-field region. The observed low-field NMR is highly temperature-dependent. It flattens rapidly as increases and almost disappears at K. Such temperature-sensitive NMR is typically attributed to either the Kondo effects [21, 22] or WL [14]. The origin of the NMR will be discussed in detail in the following sections.
At high fields, the out-of-plane MR is dominated by the classical parabolic MR, which shows little variation between 0.3 K and 20 K, as presented in Figs. 2(a) and (b). Since the amplitude of classical MR is primarily determined by the concentration and mobility of charge carriers, such a behavior indicates that the carrier concentration and mobility of TiSe2 flakes do not change significantly over this temperature range.
At K, the Hall resistivity [] is almost independent of temperature, as illustrated in Figs. 3(a) and (b), further confirming that the carrier density of TiSe2 remains constant at low temperatures. The slightly curved curves suggest the presence of multiple types of charge carriers in TiSe2, consistent with the results from ARPES measurements [10].
III.3 Possible Origins of Negative Magnetoresistance
III.3.1 The Kondo Effect
The NMR caused by the Kondo effect is expected to be isotropic in magnetic fields. To reliably compare the NMR data taken under perpendicular and in-plane magnetic fields, we first subtract the background of classical MR in the curves. The classical MR is obtained by fitting to the and data simultaneously using a two-band Drude model [23, 24] consisting of an electron band and a hole band:
and
where is the elementary charge, , , , and correspond to the 2D carrier density and mobility of the electron and hole bands, respectively. Here, the electron band represents the 3D-like electron pocket at the point, while the hole band is associated with the 2D-like hole pocket at the point [23]. The fitting results are plotted as dashed lines in Figs. 3(a)-(d), and more details are available in the Supplemental Material [15].
After subtracting the background, the net resistivity change due to the out-of-plane NMR is obtained and plotted together with the data in Figs. 3(e) and (f). The procedure for the classical-background subtraction is presented in the Supplemental Material [15]. Clearly, for both devices, the curves are much broader than the curves, indicating a high degree of anisotropy in the NMR. Such anisotropy is inconsistent with the expectations of the Kondo effect, which predicts an isotropic NMR independent of field direction [21, 22]. Therefore, the possibility of the Kondo effect can be safely excluded.
III.3.2 Two-Dimensional Weak Localization
In previous work, the NMR in TiSe2 flakes was attributed to the 2D WL effect [14]. However, in our experiment, the NMR is observed under both perpendicular and in-plane magnetic fields, as shown in Fig. 2. The presence of in-plane NMR strongly points to a 3D origin, rendering the explanation of 2D WL unlikely.
To further investigate whether the observed NMR can be ascribed to the 2D WL effect, we carefully analyzed the magnetoconductance (MC) data using the 2D WL theory. In this theory, the low-field MC due to 2D WL is described by the simplified Hikami-Larkin-Nagaoka (HLN) equation [25]
where is the change in sheet conductance under magnetic fields, is the phase coherence length, is the digamma function, and is a prefactor determined by the number of independent diffusive transport channels.
To check whether the observed NMR can be attributed to 2D WL, we fitted the HLN equation to the low-field MC data, as plotted in Fig 4. The contribution from classical MR had been removed prior to the fitting using the method described in Sec. III.3.1. Although the fitting results look reasonably good, the extracted parameters do not align with the theoretical expectations. First, the parameter was found to increase monotonically with decreasing , reaching in device A at K, far beyond the expected value of for 2D WL in a multi-band system with inter-band scattering. In addition, at K, the data are highly scattered and lack a clear trend, which is inconsistent with the power-law temperature dependence predicted by the 2D WL theory.
We also note that similar large values of , which far exceeds the expected value for 2D weak antilocalization (WAL), have been reported in several bulk topological semimetals [26, 27, 28]. As in our case, these observations cannot be explained by conventional 2D WAL and instead point to a 3D origin.
Based on the analysis presented above, we conclude that the NMR observed in this experiment does not originate from the 2D WL effect.
III.3.3 Anisotropic Three-Dimensional Weak Localization
After ruling out the possibility of the Kondo and 2D WL effects, the only remaining known mechanism to explain the observed NMR in TiSe2 flakes is the 3D WL effect. Unlike the 2D WL case, the theory of 3D WL predicts the occurrence of NMR in both out-of-plane and in-plane magnetic fields. For anisotropic systems with , where and are the diffusion coefficients in the out-of-plane and in-plane directions, the low-field conductivity correction due to 3D WL is given by [29, 30]
and
Here, and denote the phase coherence lengths in the perpendicular and in-plane directions, is the anisotropy parameter, and the function is expressed as
Figs. 5(a)-(d) show the 3D magnetoconductivity data for both devices. As usual, the background of classical MR has been removed. Notably, the amplitudes of the curves are very similar in the two devices, despite the distinct differences in the amplitudes of the curves shown in Figs. 4(a)-(b). This strongly implies that the observed NMR is essentially due to a quantum correction to the conductivity, rather than to the conductance, which further confirms the 3D nature of the NMR.
To determine whether the observed NMR can be well described by the 3D WL thoery, we analyzed the low-field data from devices A and B using Eqs. (4) and (5). Here, it is important to note that the anisotropy parameter appears in both Eqs. (4) and (5) and therefore should not be taken as an independent parameter when fitting to the and curves. However, in this experiment, the and curves were not measured at exactly the same temperature. Therefore, during the fitting, we aimed to collapse the data extracted from the and curves onto a common trend line to ensure consistency.
The 3D WL analysis yields consistent results. For both devices, the fitting curves agree well with the low-field magnetoconductivity data, as shown in Figs. 5(a)-(d). However, surprisingly, unlike the 3D WL effect previously reported in other materials [31, 32], the anisotropy parameter obtained in this experiment is not constant, but highly temperature-dependent. As shown in Fig. 5(e), at low temperatures, the extracted increases sharply from at K to at K.
Clearly, such a highly temperature-sensitive anisotropy parameter cannot be attributed to the spatial anisotropy of the effective mass tensor. Although the exact origin of such an anomalous temperature dependence of is not yet clear, we speculate that it may be arise from a possible temperature-dependent anisotropy in the inelastic scattering rate in TiSe2.
Fig. 5(f) shows the data extracted from the fitting. As decreases, increases monotonically, whereas shows no clear trend. At low temperatures, falls in the range of 10–20 nm, which is significantly smaller than the flake thickness of approximately 100 nm. This indicates that, in terms of diffusive transport, the TiSe2 flakes should be regarded as 3D systems, consistent with the presence of 3D WL.
The extracted values of are comparable to those typically obtained in other materials, whereas the values of are noticeably smaller. Such a pronounced anisotropy in phase coherence length may originate from both the anisotropic effective mass and scattering rates of electrons in TiSe2, although the exact mechanism remains to be clarified in future studies. If our interpretation is correct, a dimensional crossover from 3D to 2D WL should occur when the flake thickness is reduced below 20 nm. We plan to investigate this transition in future work.
The in-plane phase coherence length reaches about 100 nm at low temperatures and follows a temperature dependence of below 1 K, as shown by the dashed line in Fig. 5(f). Theoretically, the exponent in the relation is expected to take values of , 1 and 0.75 for systems in which the inelastic scattering is dominated by electron-phonon interaction, EEI in the clean limit and EEI in the dirty limit, respectively [20]. Clearly, the exponent obtained in this experiment does not match any of these theoretically predicted values. Although a similar value of has been reported in the 3D WL of nitrogen-doped ultrananocrystalline diamond films [32], the exact origin of this deviation remains unclear. Further research is needed to fully elucidate the discrepancy.
IV Conclusion
In summary, we observed an unusual anisotropic NMR in exfoliated TiSe2 nanoflakes. Unlike the NMR commonly seen in other 2D materials, our results cannot be explained by either the 2D WL effect or the Kondo effect. We tentatively attribute the observed NMR to the anisotropic 3D WL effect. Our results provide direct experimental signatures of phase-coherent transport processes in TiSe2 and may be helpful for advancing the understanding of phase-coherent transport phenomena in other layered TMD materials with similar anisotropic characteristics.
Acknowledgements.
This work is supported by the National Key Research and Development Program of China (Grant Nos. 2022YFA1403400, 2020YFA0308800); by the Beijing Natural Science Foundation (Grant No. Z210006); by the Beijing National Laboratory for Condensed Matter Physics (Grant No. 2023BNLCMPKF007).References
- [1] F. J. Di Salvo, D. E. Moncton, and J. V. Waszczak, Electronic properties and superlattice formation in the semimetal TiSe2, Phys. Rev. B 14, 4321 (1976).
- [2] K. C. Woo, F. C. Brown, W. L. McMillan, R. J. Miller, M. J. Schaffman, and M. P. Sears, Superlattice formation in titanium diselenide, Phys. Rev. B 14, 3242 (1976).
- [3] K. Rossnagel, L. Kipp, and M. Skibowski, Charge-density-wave phase transition in 1-TiSe2: Excitonic insulator versus band-type Jahn-Teller mechanism, Phys. Rev. B 65, 235101 (2002).
- [4] J. van Wezel, P. Nahai-Williamson, and S. S. Saxena, Exciton-phonon-driven charge density wave in TiSe2, Phys. Rev. B 81, 165109 (2010).
- [5] K. Rossnagel, On the origin of charge-density waves in select layered transition-metal dichalcogenides, J. Phys. Condens. Matter 23, 213001 (2011).
- [6] C. H. Chen, W. Fabian, F. C. Brown, K. C. Woo, B. Davies, B. DeLong, and A. H. Thompson, Angle-resolved photoemission studies of the band structure of TiSe2 and TiS2, Phys. Rev. B 21, 615 (1980).
- [7] O. Anderson, R. Manzke, and M. Skibowski, Three-dimensional and relativistic effects in layered -TiSe2, Phys. Rev. Lett. 55, 2188 (1985).
- [8] O. Anderson, G. Karschnick, R. Manzke, and M. Skibowski, The phase transition in the electronic structure of -TiSe2, Solid State Commun. 53, 339 (1985).
- [9] J. C. Rasch, T. Stemmler, B. Müller, L. Dudy, and R. Manzke, -TiSe2: Semimetal or semiconductor?, Phys. Rev. Lett. 101, 237602 (2008).
- [10] M. D. Watson, O. J. Clark, F. Mazzola, I. Marković, V. Sunko, T. K. Kim, K. Rossnagel, and P. D. C. King, Orbital- and -selective hybridization of Se and Ti states in the charge density wave phase of TiSe2, Phys. Rev. Lett. 122, 76404 (2019).
- [11] D. J. Campbell, C. Eckberg, P. Y. Zavalij, H. Kung, E. Razzoli, M. Michiardi, C. Jozwiak, A. Bostwick, E. Rotenberg, and A. Damascelli, Intrinsic insulating ground state in transition metal dichalcogenide TiSe2, Phys. Rev. Mater. 3, 53402 (2019).
- [12] J. M. Moya, C. L. Huang, J. Choe, G. Costin, M. S. Foster, and E. Morosan, Effect of synthesis conditions on the electrical resistivity of TiSe2, Phys. Rev. Mater. 3, 84005 (2019).
- [13] S. H. Huang, G. J. Shu, W. W. Pai, H. L. Liu, and F. C. Chou, Tunable Se vacancy defects and the unconventional charge density wave in -TiSe2-δ, Phys. Rev. B 95, 45310 (2017).
- [14] Z. Li, J. Zhang, M. Zhou, Y. Zeng, D. Qi, and J. Li, Weak localization in -TiSe2 microflakes, Phys. Rev. B 101, 155111 (2020).
- [15] See Supplemental Material for details regarding the magnetoresistance data under different currents, estimation of uncertainty in resistivity calculation and subtraction of the classical magnetoresistance background.
- [16] C. Monney, E. F. Schwier, M. G. Garnier, N. Mariotti, C. Didiot, H. Cercellier, J. Marcus, H. Berger, A. N. Titov, and H. Beck, Probing the exciton condensate phase in -TiSe2 with photoemission, New J. Phys. 12, 125019 (2010).
- [17] S. Koley, M. S. Laad, N. S. Vidhyadhiraja, and A. Taraphder, Preformed excitons, orbital selectivity, and charge density wave order in -TiSe2, Phys. Rev. B 90, 115146 (2014).
- [18] B. Hildebrand, T. Jaouen, C. Didiot, E. Razzoli, G. Monney, M. Mottas, A. Ubaldini, H. Berger, C. Barreteau, and H. Beck, Short-range phase coherence and origin of the -TiSe2 charge density wave, Phys. Rev. B 93, 125140 (2016).
- [19] M. D. Watson, A. M. Beales, and P. D. King, On the origin of the anomalous peak in the resistivity of TiSe2, Phys. Rev. B 99, 195142 (2019).
- [20] P. A. Lee and T. V. Ramakrishnan, Disordered electronic systems, Rev. Mod. Phys. 57, 287 (1985).
- [21] S. Barua, M. C. Hatnean, M. R. Lees, and G. Balakrishnan, Signatures of the Kondo effect in VSe2, Sci. Rep. 7, 1 (2017).
- [22] H. Liu, Y. Xue, J. Shi, R. A. Guzman, P. Zhang, Z. Zhou, Y. He, C. Bian, L. Wu, and R. Ma, Observation of the Kondo effect in multilayer single-crystalline VTe2 nanoplates, Nano. Lett. 19, 8572 (2019).
- [23] P. Knowles, B. Yang, T. Muramatsu, O. Moulding, J. Buhot, C. J. Sayers, E. Da Como, and S. Friedemann, Fermi surface reconstruction and electron dynamics at the charge-density-wave transition in TiSe2, Phys. Rev. Lett. 124, 167602 (2020).
- [24] X. Luo, C. Fang, C. Wan, J. Cai, Y. Liu, X. Han, Z. Lu, W. Shi, R. Xiong, and Z. Zeng, Magnetoresistance and Hall resistivity of semimetal WTe2 ultrathin flakes, Nanotechnology 28, 145704 (2017).
- [25] S. Hikami, A. I. Larkin, and Y. Nagaoka, Spin-orbit interaction and magnetoresistance in the two dimensional random system, Prog. Theor. Phys. 63, 707 (1980).
- [26] Z. Hou, Y. Wang, G. Xu, X. Zhang, E. Liu, W. Wang, Z. Liu, X. Xi, W. Wang, and G. Wu, Transition from semiconducting to metallic-like conducting and weak antilocalization effect in single crystals of LuPtSb, Appl Phys Lett 106, 102102 (2015).
- [27] A. Laha, S. Malick, R. Singha, P. Mandal, P. Rambabu, V. Kanchana, and Z. Hossain, Magnetotransport properties of the correlated topological nodal-line semimetal YbCdGe, Phys. Rev. B 99, 241102 (2019).
- [28] S. Roy, B. Das, A. Ghosh, A. Bera, C. K. Barman, M. Mondal, A. Alam, and P. Mandal, Chiral anomaly and weak antilocalization effects in the Weyl semimetal EuAuSb, Phys. Rev. B 110, 85145 (2024).
- [29] A. Kawabata, Theory of negative magnetoresistance I. Application to heavily doped semiconductors, J. Phys. Soc. Jpn. 49, 628 (1980).
- [30] W. Szott, C. Jedrzejek, and W. P. Kirk, Theory of weak localization in a superlattice, Phys. Rev. B 40, 1790 (1989).
- [31] Y. A. Pusep, M. B. Ribeiro, H. Arakaki, C. A. De Souza, P. A. Zanello, A. J. Chiquito, S. Malzer, and G. H. Döhler, Anisotropy of quantum interference in disordered GaAs/AlxGa1-xAs superlattices, Phys. Rev. B 68, 195207 (2003).
- [32] K. V. Shah, D. Churochkin, Z. Chiguvare, and S. Bhattacharyya, Anisotropic weakly localized transport in nitrogen-doped ultrananocrystalline diamond films, Phys. Rev. B 82, 184206 (2010).
Supplemental Material
Appendix S1 Comparison of Magnetoresistance Data under Different Currents
To rule out the possibility of overheating caused by the excitation current, we compared the magnetoresistance data measured at base temperature using three different excitation currents: 1 A, 3 A and 5 A. As shown in Fig. S1(a), the magnetoresistance curves for 1 A and 3 A are closely matched, whereas the curve for 5 A appears noticeably flatter. These results indicate that a 3 A excitation current does not cause overheating in our TiSe2 samples. Additionally, the noise level at 3 A is significantly lower than that at 1 A. Therefore, we select a 3 A excitation current for our main measurements.
Appendix S2 Estimation of Uncertainty in Resistivity Calculation
The length-to-width ratio of the devices was calculated using , where is the distance between the midpoints of the two voltage (potential) contacts, and is the width of the TiSe2 flake, as illustrated in Figs. S1(b) and (c). The selected TiSe2 flakes have well-defined geometries, which minimize the uncertainty in estimating the length-to-width ratio due to sample shape. Although the presence of voltage contacts may perturb the current distribution and introduce some uncertainty in determining the effective length, this uncertainty can be reasonably estimated by taking and as the upper and lower bounds for the length, respectively, as indicated in Figs. S1(b) and (c). For both devices, the maximum relative uncertainty in the length-to-width ratio is approximately , which would correspondingly result in an uncertainty in the calculated resistivity , as well as in the prefactor . However, this level of uncertainty does not affect the validity of our main conclusions.
Appendix S3 Fitting Details of the Classical Magnetoresistance Background
To reliably obatin the classical parabolic background in the curves, we fitted the two-band Drude model simultaneously to the high-field ( T) and data. According to the two-band Drude model, the longitudinal and transverse resistivity are respectively given by
and
where , , and are carrier density and mobility of the two bands. Note that here , are sign-carrying variables, where a positive value indicates p-type carriers and a negative value indicates n-type carriers.
| ( cm3) | ( cm3) | (cm2V-1s-1) | (cm2V-1s-1) | |
| Fitted values | ||||
| Upper bounds | ||||
| Lower bounds |
Figure S2 illustrates the two-band fitting of the high-field and curves measured at base temperature. To estimate the uncertainty range of the fitting parameters, we vary around its optimal value and then optimize the other parameters to minimize the residuals (defined as ) between the measured and fitted curves. We then compare the standard deviation of the residuals with its minimized value obtained under the best-fit condition. The boundary of the parameter range is defined as the point at which the standard deviation of the residuals reaches ten times its best-fit value. The fitted values and uncertainty ranges of all fitting parameters are listed in Table S1.
Appendix S4 Procedure for Classical Magnetoresistance Background Subtraction
As described in the main text, the raw data contain a large positive classical magnetoresistance (MR) in addition to a smaller quantum correction. To isolate the quantum contribution, we fitted the two-band Drude model to the high-field part of the data, thereby obtaining the corresponding classical background . The preliminary background-subtracted curve is then defined as
Since this subtraction procedure uses the high-field region as the reference point, does not naturally vanish at . To make the result directly comparable with theoretical descriptions of quantum corrections (such as the Kondo effect), which characterize the field-induced relative change of resistivity from its zero-field value, we further shift the curve by subtracting :
This procedure ensures that , consistent with the theoretical framework. The resulting curves shown in Figs. 3(e)–(f) of the main text therefore represent the relative change of resistivity under magnetic field after removing the classical background. For illustration, Fig. S3 shows an example of this background subtraction process using the data from Device A.